OpenStudy (joehoffy):

Check my answers How much more would the value of y be in the table than its value on the graph when x = 11? 110 150 215 275

OpenStudy (joehoffy):

http://prntscr.com/a5gj0g

OpenStudy (cresciez):

http://openstudy.com/updates/54dce95ee4b0ffc24cc5f26b Check this out :) It has the answer and it'll help walk you through the problem <3

OpenStudy (joehoffy):

i think its B

rebeccaxhawaii (rebeccaxhawaii):

well 180/3=60 240/4= 60

rebeccaxhawaii (rebeccaxhawaii):

so it would be 11 x 60

rebeccaxhawaii (rebeccaxhawaii):

im not seeing the answer though

OpenStudy (joehoffy):

is it B?

rebeccaxhawaii (rebeccaxhawaii):

@AloneS why you leave me hanging though you were my OG smh

alones (alones):

OG? :3

rebeccaxhawaii (rebeccaxhawaii):

its a reference

alones (alones):

Okay so what did you get?

OpenStudy (goldphenoix):

In both diagram, they are linear. In the left chart, everytime x increases by 1, y increases by 60. In the right chart, everytime x increases by 1, y increases by 35. Based on my interpretation of the question, on the right diagram, if x = 11, then y = 35*11 = 385. On the left diagram, if x = 11, y = 60*11 = 660. Finally, subtract to get 660-385= 275

rebeccaxhawaii (rebeccaxhawaii):

ohhhhh i didnt look at the graph correctly i thought they were the same

rebeccaxhawaii (rebeccaxhawaii):

but yes @GoldPhenoix is right

alones (alones):

mhm -.- @GoldPhenoix thank you..i guess i didn't had to rewrite it all xP